How long will it take $700 to triple itself at 10% simple interest?
Answer by RAY100(1637)

$700*3=$2100
gain is 2100-700=1400 which is interest
remember I=PRT
1400=700*.10*T
2=.10 *T
T=20 yrs
check
A=P+Prt
2100=700 +700*.10*20=2100,,,ok